OpenStudy (anonymous):

isolated metal sphere has radius r. when charged to potential v the charge on the sphere is q state an expression in trems of r q for potential V of sphere.

OpenStudy (experimentx):

google for capacitance of isolated sphere.

OpenStudy (anonymous):

I dont understand at that .will you help me ?

OpenStudy (anonymous):



OpenStudy (anonymous):

Initial charge does does not matter. Potential of metal sphere at any instant is given by V=k*q/r -------->k=1/(4*pi*epsilon_naught) =>q = V*r/k